OpenStudy (anonymous):

Solve. 1 over 25 = 5^x + 4

OpenStudy (anonymous):

1/25 = 5^(x + 4) I'm assuming. 1/25 = 5^-2, so... 5^-2 = 5^(x + 4) Log base 5 both sides. -2 = x + 4 x = -6